A rod of uniform cross section area A and length L has a weight W. It is suspended vertically from a fixed support. The material  of the rod is homogeneous and the modulud of elasticity is Y. If a concentrated force F is also applied at the free end of the rod, then find the total elongation of the rod. 

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Sanchit Gupta

Ans: First of all let's calculate the total extension of the rod.

We know that Stress/Strain =Υ. As the mass is uniformly distributed and hence it is acting at a distance L/2 from the top

so, W/A=ΥΔl/(L/2)

Hence Δl=WL/2AY;

now the external force F can act as both as a Tensile as well as compressive force.

hence the total change in length will be

case 1 Δl=(WL+F)L/2AY When F is tensile;

case 2Δl=(WL-F)L/2AY when F is compressive.

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Sanchit Gupta !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur